OpenStudy (anonymous):

Can someone help me on this one? What is the momentum of a 76-kilogram horse running at a speed of 45 kilometers/hour? A) 2.5 x 10^2 kilograms*meters/seconds B) 3.4 x 10^4 kilograms*meters/seconds C) 3.4 x 10^3 kilograms*meters/seconds D) 9.5 x 10^2 kilograms*meters/seconds

OpenStudy (anonymous):

momentum = mass * velocity or momentum = kilogram of horse * kilometers/hr does that help.. or make it worse? haha ^_^

OpenStudy (anonymous):

So it's B or C right? Not sure which one or if it actually is one of those.

OpenStudy (anonymous):

ok ok, I'm getting an idea where you are at. soooo, anytime... Everytime the momentum equation is used, the velocity, or speed, needs to be in meters/second. and the question gave us kilometers per hour, sooo we need to convert the speed of 45kilometers/hour into _?_meters/second. can you do that? if not I can show you how ^_^

OpenStudy (anonymous):

@FearTheGrace

OpenStudy (anonymous):

I got it wrong because I was running out of time and just chose a random one haha, thank you though.

OpenStudy (anonymous):

dang :( for next time here it is:\[\frac{ 45 kilometers }{ hour } \times \frac{ 1 hour }{ 60 minutes}\times \frac{ 1minute }{ 60seconds }\times \frac{ 1000meter }{ 1kilometer} = \frac{ 12.5meters }{ second }\] 12.5*76 = [answer]

OpenStudy (anonymous):

Well thank you, you're so awesome haha.

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

What is the momentum of a 76-kilogram horse running at a speed of 45 kilometers/hour?